# DATransNet: Dynamic Attention Transformer Network for Infrared Small Target Detection

### v3

@@ III-C3 Studies of GFEM
As mentioned in III, we conduct the experiments on the modules in the GFEM. As we can see, the network combining the two modules performs better.
@@ III-D Comparsion with State-of-the-art (SOTA) Methods
To evaluate the effectiveness of our proposed model, we compare it with several SOTA methods on NUDT-SIRST and IRSTD-1K datasets, including SCTransNet [16], UIUNet [8], ACM [6], ALCNet [14], AGPCNet [15], RDIAN [13], and DNANet [7]. As shown in Tab. IV, and our method performs excellently. We utilize DATransNet with dilation ratios of 1 and 5 as the baseline models for comparative evaluation. The comparison experiments include qualitative results and quantitative results, receptively.
III-D1 Qualitative Result
As shown in Fig. 5, our model preserves the shape of targets more closely to ground truth with low missed detections (green boxes) and false alarms (yellow boxes) compared to other methods, achieve satisfied results.
Module Metrics NUDT-SIRST IRSTD-1K
Params(M) mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$ mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$
ACM[6] 0.40 70.97 82.99 97.67 7.33 63.78 77.91 93.93 5.81
RDIAN[13] 0.90 87.78 93.47 97.88 9.67 64.08 78.12 93.34 5.30
ALCNet[14] 0.37 92.45 96.08 98.94 2.62 65.10 78.88 93.26 7.90
AGPCNet[15] 13.36 86.43 92.72 98.20 3.70 67.54 80.64 94.27 4.76
DNANet[7] 4.69 93.73 96.70 99.25 4.55 65.79 79.23 93.60 5.96
UIU-Net[8] 50.54 94.11 96.96 97.98 0.74 68.06 80.91 92.59 5.85
SCTransNet[16] 11.19 93.83 96.77 98.20 1.56 67.97 80.90 93.26 6.24
Ours 4.46 94.93 97.39 99.04 2.00 68.26 81.11 95.95 4.42
@@ III-D2 Quantitative Results
As shown in the Tab. IV, our method achieves impressive performance, including an $mIoU$ of $94.93\%$, an $F1$ score of $97.39\%$, and a $Pd$ of $99.04\%$, a $Pa$ of $2.00\times 10^{-6}$. Similarly, on the IRSTD-1K dataset, our method is a leading network.
Figure 6: ROC curves of varied networks on NUDT-SIRST datasets.
Furthermore, we used the ROC curve to analyze the performance of various models, as depicted in Fig. 6. Although UIU-Net achieves commendable results in $mIoU$, its AUC scores are comparatively weak. In contrast, ALCNet performance shows a significant divergence, with subpar outcomes in $mIoU$ but relatively strong performance in ROC metrics. The ROC curve corroborates the findings in Tab. IV. The model proposed in this article consistently shows good performance in both ROC and $mIoU$.
@@ IV Conclusion
In this letter, we propose an ISTD network that enhances detection performance. The DATrans was proposed to enhance the local gradient feature extraction and detection performance of the network. Additionally, we propose the GFEM, which focuses on global perception across the entire feature map and learns relationships between distant pixels, preventing DATrans from overlooking background information due to its emphasis on local gradient differences. Then the local and global feature are fused. We conducted extensive experiments indicating satisfied results with fewer parameters. The model is a back-box approach because its capability for dynamic weight adjustments lacks mathematical proof. Moreover, it suffers from false alarms when processing high-noise images. To address these limitations, we aim to improve the model in two aspects. At first, we plan to improve the interpretability of the network. Then, we would like to integrate temporal information into the model to improve its robustness.

### v4

@@ III-C3 Studies of GFEM
As mentioned in III, we conduct the experiments on the modules in the GFEM. The network combining the two modules performs better with a slight increase in computational complexity.
@@ III-D Comparsion with State-of-the-art (SOTA) Methods
We compare it with several SOTA methods on NUDT-SIRST and IRSTD-1K, including SCTransNet [13], UIUNet [8], ACM [6], ALCNet [11], AGPCNet [12], RDIAN [10], and DNANet [7]. In addition, deep supervision is not employed to ensure equality in any of the networks. We used DATransNet with dilation ratios of 1 and 3 for comparative evaluation. The results of the experiments include qualitative and quantitative results.
III-D1 Qualitative Results
As shown in Fig. 6, our model preserves the shape of targets more closely to ground truth with low missed detections and false alarms.
@@ III-D2 Quantitative Results
As shown in the Tab. IV, our method achieves impressive performance, including a $mIoU$ of $94.93\%$, an $F1$ score of $97.39\%$, and a $Pd$ of $99.04\%$, a $F_{a}$ of $2.00\times 10^{-6}$. Similarly, on the IRSTD-1K dataset, our method is a leading network.
Figure 7: ROC curves of varied networks on NUDT-SIRST and IRSTD-1K datasets.
Furthermore, we used the ROC curve to analyze the performance of various models, as depicted in Fig. 7. The model proposed in this article consistently shows good performance in both ROC and $mIoU$.
@@ IV Conclusion
In this Letter, we propose an ISTD network that enhances detection performance. DATrans was proposed to enhance the network’s local gradient feature extraction and detection performance. Additionally, we propose the GFEM, which focuses on global perception across the entire feature map and learns relationships between distant pixels. We have conducted extensive experiments, indicating satisfied results with fewer parameters. The model is still a black-box approach because its capability for dynamic weight adjustments lacks mathematical proof. Moreover, it suffers from false alarms when processing high-noise images. To address these limitations, we aim to improve the model in two ways. At first, we plan to improve the interpretability of the network. Then, we would like to integrate temporal information into the model to improve its robustness.